1-Your mother just inherited $500,000. If she invests the money in a well-diversified, low-cost mutual fund returning 8% per year, how many years will it take her investment to become worth $1,000,000?

2- You just graduated from college and landed your first "real" job, which pays $70,000 a year. In 19 years how much will you need to earn to maintain the same purchasing power if inflation is 3.00% per year?
Answer #1

Present value = Future value/(1+i)^n

i = interest rate per period

n= number of periods

1)

500000 = 1000000/(1+8%)^n

=>

n = ln(1000000/500000)/ln(1.08)

= 9.01 years

= 9 years

2)

future value = 70000 * (1+3%)^19

= 122745.42